The sum of two numbers is 28. the difference of those numbers is 38, what is the second number
Answer by richwmiller(17219)

a+b=28
a-b=38
2a=66
a=33
33+b=28
b=-5
and
a+b=288
a-b=38
2a=326
a=163
288-163=125
b=125